Find the 20th term of the arithmetic sequence in which A6=70 and A9=91.
One of the following is the correct answer. Which one?
A) 1061
B) 263
C) 175
D) 168
E) 133

Answer by greenestamps(13200)



A9 is A6 plus the common difference 9-6=3 times:

91=70+3d
21=3d
d=7

A20 is A9 plus the common difference 20-9=11 times:

91+11(7)=91+77=168

ANSWER D)